# X-Vine Models for Multivariate Extremes

### v1

@@ 3 Sklar’s theorem for tail copula densities
Let $r:(0,\infty)^{d}\to[0,\infty)$ be a $d$-variate tail copula density, that is, properties (2.4) and (2.5) hold. Recall its multivariate margins in (2.7). For non-empty, disjoint subsets $I,J\subset[d]$ and for $\bm{x}_{I}\in(0,\infty)^{I}$ and $\bm{x}_{J}\in(0,\infty)^{J}$ such that $r_{J}(\bm{x}_{J})>0$, define
$\forall\bm{x}_{I\cup J}\in(0,\infty)^{I\cup J},\qquad r_{I|J}(\bm{x}_{I}|\bm{x}_{J}):=\frac{r_{I\cup J}(\bm{x}_{I\cup J})}{r_{J}(\bm{x}_{J})}.$
Viewing the “conditional” tail copula density $r_{I|J}$ as an $|I|$-variate probability density, we can decompose it into $|I|$ univariate probability densities and a copula density.
@@
Let $r$ be a $d$-variate tail copula density and let $I,J\subset[d]$ be non-empty and disjoint. For $\bm{x}_{J}\in(0,\infty)^{J}$ such that $r_{J}(\bm{x}_{J})>0$, the function
$r_{I|J}(\,\cdot\,|\bm{x}_{J}):(0,\infty)^{I}\to\left[0,\infty\right),\quad\bm{x}_{I}\mapsto r_{I|J}(\bm{x}_{I}|\bm{x}_{J}),$
is a probability density on $(0,\infty)^{I}$. For $i\in I$, it has marginal probability density $x_{i}\mapsto r_{i|J}(x_{i}|\bm{x}_{J})$ and cumulative distribution function
$x_{i}\mapsto R_{i|J}(x_{i}|\bm{x}_{J}):=\int_{0}^{x_{i}}r_{i|J}(t|\bm{x}_{J})\,\mathrm{d}t,$
with quantile function $u_{i}\mapsto R_{i|J}^{-1}(u_{i}|\bm{x}_{J})$. If $|I|\geqslant 2$, then the copula density of $r_{I|J}(\,\cdot\,|\bm{x}_{J})$ is
[…]
@@
In Proposition 3.1, for $\bm{x}_{J}\in(0,\infty)^{J}$ such that $r_{J}(\bm{x}_{J})>0$, the family $\left\{{r_{I|J}(\,\cdot\,|t\bm{x}_{J}):t\in(0,\infty)}\right\}$ of probability densities is a scale family: if the random vector $\bm{\xi}_{I}$ has probability density $r_{I|J}(\,\cdot\,|\bm{x}_{J})$ then the random vector $t\bm{\xi}_{I}$ has probability density $r_{I|J}(\,\cdot\,|t\bm{x}_{J})$. As a consequence, for all $t\in(0,\infty)$, $\bm{x}_{I}\in(0,\infty)^{I}$, $\bm{u}_{I}\in(0,\infty)^{I}$ and $i\in I$, we have
$r_{i|J}(tx_{i}|t\bm{x}_{J})$ $=r_{i|J}(x_{i}|\bm{x}_{J}),\quad r_{i|J}^{-1}(u_{i}|t\bm{x}_{J})=t\cdot r_{i|J}^{-1}(u_{i}|\bm{x}_{J}),\quad\text{ and}$
@@
Let $\bm{Y}$ be the multivariate Pareto vector in (2.11), suppose $d\geqslant 3$, and let $i,j$ be distinct elements in $[d]$, with $J=[d]\setminus\left\{{i,j}\right\}$. Provided $r$ is positive and continuous, $Y_{i}$ and $Y_{j}$ are conditionally independent given $\bm{Y}_{J}$ in the sense of Definition 1 in Engelke and Hitz (2020) if and only if $c_{i,j;J}(u_{i},u_{j};\bm{x}_{J})=1$ for all $(u_{i},u_{j})\in(0,1)^{2}$ and $\bm{x}_{J}\in(0,\infty)^{J}$, the density of the independence copula. This statement follows by combining the factorization property in Proposition 1 in Engelke and Hitz (2020) with equations (2.12) and (3.1); see also (D.1) in Appendix D for an expression of $c_{I;J}$ directly in terms of the exponent measure density $\lambda$.
@@
Sklar’s theorem for tail copula densities (Proposition 3.1) admits a converse part, stating how to combine several tail copula densities along a scale-invariant family of copula densities to form a tail copula density in a higher dimension. Proposition C.1 in Appendix C gives a precise statement. In the remainder of the paper, we will not use this result as such, although the X-vine construction in Section 5 is in the same spirit.
@@ 4.1 Scaled extremal Dirichlet model
In dimension $d\geqslant 2$, consider parameters $\alpha_{1},\ldots,\alpha_{d}>0$ and $\rho>-\min(\alpha_{1},\ldots,\alpha_{d})$ with $\rho\neq 0$. The angular measure density of the scaled extremal Dirichlet model is introduced in Belzile and Nešlehová (2017, Proposition 5). By Eq. (2.12) and Lemma B.1, the corresponding tail copula density is
@@
In McNeil and Nešlehová (2010, Definition 3), the family of Liouville copulas is defined as the collection of survival copulas of random vectors of the form $\bm{Y}=S\bm{W}$, where $\bm{W}$ has a Dirichlet distribution on the unit simplex and is independent of the positive random variable $S$, whose distribution is referred to as the radial distribution. The special case where $\bm{W}$ is uniformly distributed on the unit simplex yields the family of Archimedean copulas (McNeil and Nešlehová, 2009).
@@
Let $r$ be the scaled extremal Dirichlet tail copula density in (4.1) in dimension $d\geqslant 3$. For every pair of disjoint, non-empty subsets $I,J\subset[d]$ with $|I|\geqslant 2$, the copula densities $c_{I;J}(\,\cdot\,;\bm{x}_{J})$ in (3.1) satisfy the simplifying assumption (Definition 3.4). If $\rho>0$, then $c_{I;J}$ is equal to density the $|I|$-variate Liouville copula with Dirichlet parameters $(\alpha_{i})_{i\in I}$ and with radial density on $(0,\infty)$ proportional to
$s\mapsto s^{\sum_{i\in I}\alpha_{i}-1}\cdot\left({s+1}\right)^{-\rho-\sum_{k\in I\cup J}\alpha_{k}}.$
If $\rho<0$, then $c_{I;J}$ is equal to the density of the survival copula of the above Liouville copula.
In special cases, the copula densities $c_{I;J}$ can be identified more explicitly (calculations in Appendix B.3):
@@ 4.2 Hüsler–Reiss model
Let $\Gamma$ be a $d\times d$ variogram with $d\geqslant 3$ and let $r$ be the Hüsler–Reiss tail copula density in (4.3). For every pair of disjoint, non-empty subsets $I,J\subset[d]$ with $|I|\geqslant 2$, the copula densities $c_{I;J}(\,\cdot\,;\bm{x}_{J})$ in (3.1) satisfy the simplifying assumption (Definition 3.4). The common copula density $c_{I;J}$ is equal to the $|I|$-variate Gaussian copula density (4.4) with correlation matrix $\bm{R}_{I|J}=(\Delta_{I|J}^{(k)})^{-1/2}\,\Sigma^{(k)}_{I|J}\,(\Delta_{I|J}^{(k)})^{-1/2}$ for any $k\in J$, where $\Delta_{I|J}^{(k)}$ is the diagonal matrix with the same diagonal as
$\Sigma^{(k)}_{I|J}=\Sigma^{(k)}_{II}-\Sigma^{(k)}_{IJ}\left({\Sigma^{(k)}_{JJ}}\right)^{-1}\Sigma^{(k)}_{JI}.$
The expression of the correlation matrix $\bm{R}_{I|J}$ in Proposition 4.2 depends on the choice of $k\in J$, although the actual matrix does not. This is because the proof is based on identity (4.3) for the tail copula density. In Hentschel et al. (2022), a formula for the exponent measure density is derived that does not require the choice of such an index $k$. The formula involves what is called the Hüsler–Reiss precision matrix, denoted by $\Theta$. It remains to be investigated how to express the correlation matrix $\bm{R}_{I|J}$ in terms of $\Theta$.
@@ 5 Vine decompositions of tail copula densities
In this section, we show that any tail copula density $r$ in dimension $d\geqslant 3$ can be decomposed along any regular vine sequence (Definition 2.2) on $d$ elements into $d-1$ bivariate tail copula densities and $\sum_{i=1}^{d-2}i=\binom{d-1}{2}$ bivariate copula densities. Moreover, the copula densities depending on a single conditioning variable satisfy the simplifying assumption (Definition 3.4). Conversely, starting from any regular vine sequence on $d\geqslant 3$ elements, any collection of $d-1$ bivariate tail copula densities and any collection of $\binom{d-1}{2}$ bivariate copula densities, a $d$-variate tail copula density can be assembled. We coin tail copula densities constructed in this way as X-vines. By exploiting connection (2.12) between tail copula densities and exponent measure densities, all results carry over to the latter too. For the reader’s convenience, the main formulas are spelled out in terms of exponent measure densities in Appendix D.
@@
$c_{a_{e},b_{e};D_{e}}\left({u_{a_{e}},u_{b_{e}};\bm{x}_{D_{e}}}\right)=\frac{r_{a_{e},b_{e}|D_{e}}(x_{a_{e}},x_{b_{e}}|\bm{x}_{D_{e}})}{r_{a_{e}|D_{e}}(x_{a_{e}}|\bm{x}_{D_{e}})\cdot r_{b_{e}|D_{e}}(x_{b_{e}}|\bm{x}_{D_{e}})},$
@@
Remark 5.2.
The decomposition (5.1) is based on the vine telescoping product formulas (2.13)–(2.14) in Lemma 2.4. By applying the more general formula (2.15) in Remark 2.5, we find, for every edge $f\in E_{j}$ with $j\in\left\{{1,\ldots,d-1}\right\}$, a decomposition of the marginal tail copula density on the complete union $J=A_{f}\subseteq[d]$, with cardinality $|J|=j+1$: for $\bm{x}_{J}\in(0,\infty)^{J}$, we have
$r_{J}(\bm{x}_{J})=\prod_{\begin{subarray}{c}e\in E_{1}\\ A_{e}\subseteq J\end{subarray}}r_{a_{e},b_{e}}(x_{a_{e}},x_{b_{e}})\cdot\prod_{i=2}^{|J|-1}\prod_{\begin{subarray}{c}e\in E_{i}\\ A_{e}\subseteq J\end{subarray}}c_{a_{e},b_{e};D_{e}}\left({R_{a_{e}|D_{e}}(x_{a_{e}}|\bm{x}_{D_{e}}),R_{b_{e}|D_{e}}(x_{b_{e}}|\bm{x}_{D_{e}});\bm{x}_{D_{e}}}\right).$ (5.2)
The decomposition (5.2) is true only for subsets $J\subseteq[d]$ of the form $J=A_{f}$ for some edge $f$ in the regular vine sequence.
For the pair copula $C_{a_{e},b_{e};D_{e}}(\,\cdot\,,\,\cdot\,;\bm{x}_{D_{e}})$ associated to the edge $e=(a_{e},b_{e};D_{e})$ in the regular vine sequence, consider the first-order partial derivatives
@@
A $d$-variate tail copula density $r$ is an X-vine along a regular vine sequence $\mathcal{V}$ if for each edge $e\in E_{2}\cup\cdots\cup E_{d-1}$, the pair copula densities $c_{a_{e},b_{e};D_{e}}(\,\cdot\,,\,\cdot\,;\bm{x}_{D_{e}})$ do not depend on the value of $\bm{x}_{D_{e}}\in(0,\infty)^{D_{e}}$. The $d$-variate exponent measure density $\lambda$ in (2.12) is called an X-vine if the tail copula density $r$ is an X-vine.
@@
By Lemma 3.2, a trivariate tail copula density $r$ is always an X-vine, and this along any of the three possible regular vine sequences on $\left\{{1,2,3}\right\}$. Indeed, for the vine sequence determined by $E_{1}=\left\{{\left\{{1,2}\right\},\left\{{2,3}\right\}}\right\}$, we can for instance write
$r(x_{1},x_{2},x_{3})=r_{12}(x_{1},x_{2})\,r_{23}(x_{2},x_{3})\cdot c_{13;2}\left({R_{1|2}(x_{1}|x_{2}),\,R_{3|2}(x_{3}|x_{2})}\right),$
featuring the copula density
$c_{13;2}(u_{1},u_{3})=\frac{r(x_{1},1,x_{3})}{r_{12}(x_{1},1)\,r_{23}(1,x_{3})}\quad\text{with}\quad x_{j}=R_{j|2}^{-1}(u_{j}|1),\quad j\in\left\{{1,3}\right\}.$
The function $r$ is thus completely specified by the two bivariate tail copula densities $r_{12}$ and $r_{23}$ and the bivariate copula density $c_{13;2}$. In Theorem 5.7 below, this construction principle is lifted to higher dimensions.
@@
Fig. 2 in Section 8 shows an example of an X-vine specification involving a mix of parametric models for the bivariate (tail) copula densities.
@@
is a $d$-variate tail copula density and its multivariate margins are the ones implied by the notation. In particular, $r$ is an X-vine.
Remark 5.8 (Link with regular vine copulas).
X-vine tail copula densities arise as the lower tail dependence limits of regular vine copula densities provided the pair copula densities at the edges of the first tree have the corresponding bivariate tail copula densities as lower tail dependence limits. In the passage to the limit, the regular vine sequence is preserved and so are the pair copulas at all trees starting from the second one. Max-stable dependence structures based on X-vines thus arise naturally from the max-domain of attraction condition of regular vine copulas, and the X-vine limit is particularly easy to calculate. These properties are worked out in Appendix E.
If the bivariate copula density $c_{a_{e},b_{e};D_{e}}$ associated with an edge $e\in E_{2}\cup\cdots\cup E_{d-1}$ is equal to the independence one, $c_{a_{e},b_{e};D_{e}}\equiv 1$, the corresponding factor drops out in (5.1) and the recursive formulas in (5.4) and (5.6) simplify to
$R_{a_{e}|D_{e}\cup b_{e}}\left({x_{a_{e}}|\bm{x}_{D_{e}\cup b_{e}}}\right)$ $=R_{a_{e}|D_{e}}\left({x_{a_{e}}|\bm{x}_{D_{e}}}\right),$
$R_{b_{e}|D_{e}\cup a_{e}}\left({x_{b_{e}}|\bm{x}_{D_{e}\cup a_{e}}}\right)$ $=R_{b_{e}|D_{e}}\left({x_{b_{e}}|\bm{x}_{D_{e}}}\right).$
Sparse X-vine specifications arise when $c_{a_{e},b_{e};D_{e}}\equiv 1$ for many edges $e\in E_{2}\cup\cdots\cup E_{d-1}$. Model selection of sparse vine copulas in high dimensions has been investigated in Müller and Czado (2019) and Nagler et al. (2019). The case where all pair copulas are equal to the independence one for all edges starting from a given tree is of practical importance (Brechmann et al., 2012; Brechmann and Joe, 2015).

### v2

@@ 3 Sklar’s theorem for tail copula densities
Let $r$ be a $d$-variate tail copula density, so properties (2.4) and (2.5) hold. Recall its multivariate margins in (2.7). For non-empty, disjoint $I,J\subset[d]$ and for $\bm{x}_{I}\in(0,\infty)^{I}$ and $\bm{x}_{J}\in(0,\infty)^{J}$ such that $r_{J}(\bm{x}_{J})>0$, define $r_{I|J}(\bm{x}_{I}|\bm{x}_{J}):=r_{I\cup J}(\bm{x}_{I\cup J})/r_{J}(\bm{x}_{J})$ for all $\bm{x}_{I\cup J}\in(0,\infty)^{I\cup J}$. Viewing the “conditional” tail copula density $r_{I|J}$ as an $|I|$-variate probability density, we can decompose it into $|I|$ univariate probability densities and a copula density.
@@
Let $r$ be a $d$-variate tail copula density and let $I,J\subset[d]$ be non-empty and disjoint. For $\bm{x}_{J}\in(0,\infty)^{J}$ such that $r_{J}(\bm{x}_{J})>0$, the function $r_{I|J}(\,\cdot\,|\bm{x}_{J}):(0,\infty)^{I}\to\left[0,\infty\right)$, defined via $\bm{x}_{I}\mapsto r_{I|J}(\bm{x}_{I}|\bm{x}_{J})$, is a probability density on $(0,\infty)^{I}$. For $i\in I$, it has marginal probability density $x_{i}\mapsto r_{i|J}(x_{i}|\bm{x}_{J})$ and cumulative distribution function $x_{i}\mapsto R_{i|J}(x_{i}|\bm{x}_{J}):=\int_{0}^{x_{i}}r_{i|J}(t|\bm{x}_{J})\,\mathrm{d}t$, with quantile function $u_{i}\mapsto R_{i|J}^{-1}(u_{i}|\bm{x}_{J})$. If $|I|\geqslant 2$, then the copula density of $r_{I|J}(\,\cdot\,|\bm{x}_{J})$ is
$\forall\bm{u}_{I}\in(0,\infty)^{I},\qquad c_{I;J}(\bm{u}_{I};\bm{x}_{J}):=\frac{r_{I|J}(\bm{x}_{I}|\bm{x}_{J})}{\prod_{i\in I}r_{i|J}(x_{i}|\bm{x}_{J})}\quad\text{with}\quad x_{i}=R_{i|J}^{-1}(u_{i}|\bm{x}_{J}).$ (3.1)
@@
In Proposition 3.1, for $\bm{x}_{J}\in(0,\infty)^{J}$ such that $r_{J}(\bm{x}_{J})>0$, the family $\left\{{r_{I|J}(\,\cdot\,|t\bm{x}_{J}):t\in(0,\infty)}\right\}$ of probability densities is a scale family: if the random vector $\bm{\xi}_{I}$ has probability density $r_{I|J}(\,\cdot\,|\bm{x}_{J})$, then the random vector $t\bm{\xi}_{I}$ has probability density $r_{I|J}(\,\cdot\,|t\bm{x}_{J})$. As a consequence, for all $t\in(0,\infty)$, $\bm{x}_{I}\in(0,\infty)^{I}$, $\bm{u}_{I}\in(0,1)^{I}$ and $i\in I$, we have $r_{i|J}(tx_{i}|t\bm{x}_{J})=r_{i|J}(x_{i}|\bm{x}_{J})$, $R_{i|J}^{-1}(u_{i}|t\bm{x}_{J})=t\cdot R_{i|J}^{-1}(u_{i}|\bm{x}_{J})$, and
@@
Let $\bm{Y}$ be the multivariate Pareto vector in (2.11), suppose $d\geqslant 3$, and let $i,j$ be distinct elements in $[d]$, with $J=[d]\setminus\left\{{i,j}\right\}$. Provided $r$ is positive and continuous, $Y_{i}$ and $Y_{j}$ are conditionally independent given $\bm{Y}_{J}$ in the sense of Definition 1 in Engelke and Hitz (2020) if and only if $c_{i,j;J}(u_{i},u_{j};\bm{x}_{J})=1$ for all $(u_{i},u_{j})\in(0,1)^{2}$ and $\bm{x}_{J}\in(0,\infty)^{J}$, the density of the independence copula. This statement follows by combining the factorization property in Proposition 1 in Engelke and Hitz (2020) with equations (2.12) and (3.1); see also (C.1) in Appendix C in the supplement for an expression of $c_{I;J}$ in terms of the exponent measure density $\lambda$.
@@
As a converse to Proposition 3.1, we can combine several tail copula densities along a scale-invariant family of copula densities to form a tail copula density in higher dimensions. With slight abuse of notation, write $i\cup J:=\{i\}\cup J$ for $i\in I$. A family of $k$-variate copula densities $\left\{{c(\,\cdot\,;\theta):\theta\in\Theta}\right\}$ indexed by a Borel set $\Theta\subseteq\mathbb{R}^{m}$ is jointly measurable if the map $(\bm{u},\theta)\mapsto c(\bm{u};\theta)$ is Borel measurable on $(0,1)^{k}\times\Theta$.
Proposition 3.5 (Sklar’s theorem: converse part).
$r(\bm{x}):=r_{J}(\bm{x}_{J})\cdot\prod_{i\in I}r_{i|J}(x_{i}|\bm{x}_{J})\cdot c_{I;J}\left({R_{i|J}(x_{i}|\bm{x}_{J}),i\in I;\bm{x}_{J}}\right),$ (3.3)
for $\bm{x}\in(0,\infty)^{d}$ such that $r_{J}(\bm{x}_{J})>0$ and zero otherwise, is a $d$-variate tail copula density with margins $r_{i\cup J}$ for $i\in I$.
@@ 4.1 Scaled extremal Dirichlet model
In dimension $d\geqslant 2$, consider parameters $\alpha_{1},\ldots,\alpha_{d}>0$ and $\rho>-\min(\alpha_{1},\ldots,\alpha_{d})$ with $\rho\neq 0$. The angular measure density of the scaled extremal Dirichlet model is introduced in Belzile and Nešlehová (2017, Proposition 5). By Eq. (2.12) and Lemma B.2 in the supplement, the corresponding tail copula density is
@@
In McNeil and Nešlehová (2010, Definition 3), the family of Liouville copulas is defined as the collection of survival copulas of random vectors of the form $\bm{Y}=S\bm{W}$, where $\bm{W}$ has a Dirichlet distribution on the unit simplex and is independent of the positive random variable $S$. The special case where $\bm{W}$ is uniformly distributed on the unit simplex yields the family of Archimedean copulas (McNeil and Nešlehová, 2009).
@@
Let $r$ be the scaled extremal Dirichlet tail copula density in (4.1) in dimension $d\geqslant 3$. For every pair of disjoint, non-empty subsets $I,J\subset[d]$ with $|I|\geqslant 2$, the copula densities $c_{I;J}(\,\cdot\,;\bm{x}_{J})$ in (3.1) satisfy the simplifying assumption (Definition 3.4). If $\rho>0$, then $c_{I;J}$ is equal to the density of the $|I|$-variate Liouville copula with Dirichlet parameters $(\alpha_{i})_{i\in I}$ and with radial density on $(0,\infty)$ proportional to $s\mapsto s^{\sum_{i\in I}\alpha_{i}-1}\cdot\left({s+1}\right)^{-\rho-\sum_{k\in I\cup J}\alpha_{k}}$. If $\rho<0$, then $c_{I;J}$ is equal to the density of the survival copula of the above Liouville copula.
In special cases, the copula densities $c_{I;J}$ can be identified more explicitly (calculations in Appendix B.3 in the supplement):
@@ 4.2 Hüsler–Reiss model
Let $\Gamma$ be a $d\times d$ variogram with $d\geqslant 3$ and let $r$ be the Hüsler–Reiss tail copula density in (4.3). For every pair of disjoint, non-empty subsets $I,J\subset[d]$ with $|I|\geqslant 2$, the copula densities $c_{I;J}(\,\cdot\,;\bm{x}_{J})$ in (3.1) satisfy the simplifying assumption (Definition 3.4). The copula density $c_{I;J}$ is equal to the $|I|$-variate Gaussian copula density (4.4) with correlation matrix $\bm{R}_{I|J}=(\Delta_{I|J}^{(k)})^{-1/2}\,\Sigma^{(k)}_{I|J}\,(\Delta_{I|J}^{(k)})^{-1/2}$ for any $k\in J$, where $\Delta_{I|J}^{(k)}$ is the diagonal matrix with the same diagonal as $\Sigma^{(k)}_{I|J}=\Sigma^{(k)}_{II}-\Sigma^{(k)}_{IJ}\left({\Sigma^{(k)}_{JJ}}\right)^{-1}\Sigma^{(k)}_{JI}$.
The matrix $\bm{R}_{I|J}$ in Proposition 4.2 is the correlation matrix of the conditional Gaussian distribution with covariance matrix $\Sigma^{(k)}_{I|J}$, the corresponding Schur complement. Its expression depends on the choice of $k\in J$, although the actual matrix does not. It remains to be investigated how to express it in terms of the Hüsler–Reiss precision matrix $\Theta$ introduced in Hentschel et al. (2022).
@@ 5 Vine decompositions of tail copula densities
5.1 X-vines: density decomposition and construction
In this section, we show that any tail copula density $r$ in dimension $d\geqslant 3$ can be decomposed along any regular vine sequence (Definition 2.2) on $d$ elements into $d-1$ bivariate tail copula densities and $\sum_{i=1}^{d-2}i=\binom{d-1}{2}$ bivariate copula densities. Moreover, copula densities with only a single conditioning variable satisfy the simplifying assumption (Definition 3.4). Conversely, starting from any regular vine sequence on $d\geqslant 3$ elements, any collection of $d-1$ bivariate tail copula densities and any collection of $\binom{d-1}{2}$ bivariate copula densities, a $d$-variate tail copula density can be assembled. We coin tail copula densities constructed in this way as X-vines. Appendix C in the supplement states the main results in terms of exponent measure densities.
@@ 5.1 X-vines: density decomposition and construction
$c_{a_{e},b_{e};D_{e}}\left({u_{a_{e}},u_{b_{e}};\bm{x}_{D_{e}}}\right)=\frac{r_{a_{e},b_{e}|D_{e}}(x_{a_{e}},x_{b_{e}}|\bm{x}_{D_{e}})}{r_{a_{e}|D_{e}}(x_{a_{e}}|\bm{x}_{D_{e}})\cdot r_{b_{e}|D_{e}}(x_{b_{e}}|\bm{x}_{D_{e}})},$ (5.2)
@@
The decomposition (5.1) also applies to marginal tail copula densities $r_{J}$ for index sets $J=A_{f}$ for some edge $f$ in the vine; see Remark B.7 in the supplement. For the pair copula $C_{a_{e},b_{e};D_{e}}(\,\cdot\,,\,\cdot\,;\bm{x}_{D_{e}})$ associated to the edge $e=(a_{e},b_{e};D_{e})$ in the regular vine sequence, consider the first-order partial derivatives
@@
A $d$-variate tail copula density $r$ is an X-vine along a regular vine sequence $\mathcal{V}$ if for each edge $e\in E_{2}\cup\cdots\cup E_{d-1}$, the pair copula densities $c_{a_{e},b_{e};D_{e}}(\,\cdot\,,\,\cdot\,;\bm{x}_{D_{e}})$ do not depend on the value of $\bm{x}_{D_{e}}\in(0,\infty)^{D_{e}}$.
@@
By Lemma 3.2, a trivariate tail copula density $r$ is always an X-vine, and this along any of the three possible regular vine sequences on $\left\{{1,2,3}\right\}$. For the vine determined by $E_{1}=\left\{{\left\{{1,2}\right\},\left\{{2,3}\right\}}\right\}$, for instance, we have
$r(x_{1},x_{2},x_{3})$ $=r_{12}(x_{1},x_{2})\,r_{23}(x_{2},x_{3})\cdot c_{13;2}\left({R_{1|2}(x_{1}|x_{2}),\,R_{3|2}(x_{3}|x_{2})}\right),$ (5.5)
$\text{where}\quad c_{13;2}(u_{1},u_{3})$ $=\frac{r(x_{1},1,x_{3})}{r_{12}(x_{1},1)\,r_{23}(1,x_{3})}\quad\text{with}\quad x_{j}=R_{j|2}^{-1}(u_{j}|1),\quad j\in\left\{{1,3}\right\}.$
The function $r$ is thus completely specified by the two bivariate tail copula densities $r_{12}$ and $r_{23}$ and one bivariate copula density $c_{13;2}$. The form (5.5) was already discovered for tail copula densities of D-vine copulas in Li and Wu (2013).
@@
Fig. LABEL:fig:5d_Xvine_a in Section 8 shows an example of an X-vine specification involving a mix of parametric models for the bivariate (tail) copula densities.
@@
is a $d$-variate tail copula density. For $e\in E_{1}$, the bivariate margin of $r$ is equal to $r_{e}\in\mathcal{R}$, while for $e=(a_{e},b_{e};D_{e})\in E_{2}\cup\cdots\cup E_{d-1}$, the pair copula density $c_{a_{e},b_{e};D_{e}}(\,\cdot\,,\,\cdot\,;\bm{x}_{D_{e}})$ in (5.2) is equal to $c_{a_{e},b_{e};D_{e}}\in\mathcal{C}$. In particular, $r$ is an X-vine.
In Example 5.4 with $d=3$, the vine $\mathcal{V}=(T_{1},T_{2})$ is determined by $E_{1}=\left\{{\left\{{1,2}\right\},\left\{{2,3}\right\}}\right\}$, while =$\mathcal{R}=\left\{{r_{12},r_{23}}\right\}$ and $\mathcal{C}=\left\{{c_{13;2}}\right\}$, and (5.6) reduces to (5.5).
5.2 X-vines as limits of regular vine copula densities
A natural question is whether X-vine tail copula densities arise as the lower tail dependence limits of regular vine copula densities, as introduced in Bedford and Cooke (2002). Below, we show that this is indeed the case, provided the pair copula densities at the edges of the first tree have the corresponding bivariate tail copula densities as lower tail dependence limits. In the passage to the limit, the regular vine sequence is preserved and so are the pair copulas at all trees starting from the second one.
$c(\bm{u})=\prod_{e\in E_{1}}c_{a_{e},b_{e}}(u_{a_{e}},u_{b_{e}})\cdot\prod_{j=2}^{d-1}\prod_{e\in E_{j}}c_{a_{e},b_{e};D_{e}}\left({C_{a_{e}|D_{e}}\left({u_{a_{e}}|\bm{u}_{D_{e}}}\right),\,C_{b_{e}|D_{e}}\left({u_{b_{e}}|\bm{u}_{D_{e}}}\right)}\right),$ (5.8)
$C_{a_{e},b_{e}|D_{e}}\left({u_{a_{e}},u_{b_{e}}|\bm{u}_{D_{e}}}\right)=C_{a_{e},b_{e};D_{e}}\left({C_{a_{e}|D_{e}}\left({u_{a_{e}}|\bm{u}_{D_{e}}}\right),\,C_{b_{e}|D_{e}}\left({u_{b_{e}}|\bm{u}_{D_{e}}}\right)}\right).$ (5.9)
Assumption 5.7.
Let $c$ be a $d$-variate regular vine copula density as in (5.8), with $d\geqslant 3$.
  1. (i)
    For every edge $e\in E_{1}$, there exists a bivariate tail copula density $r_{a_{e},b_{e}}$ such that $\lim_{t\searrow 0}t\cdot c_{a_{e},b_{e}}(tx,ty)=r_{a_{e},b_{e}}(x,y)$, for $(x,y)\in(0,\infty)^{2}$.
  2. (ii)
Proposition 5.8.
Under Assumption 5.7, we have $\lim_{t\searrow 0}t^{d-1}\cdot c(t\bm{x})=r(\bm{x})$ for $\bm{x}\in(0,\infty)^{d}$, where the X-vine tail copula density $r$ is generated by the triple $(\mathcal{V},\mathcal{R},\mathcal{C})$ as in Theorem 5.6 with
  • $\mathcal{V}$ the same regular vine sequence as in (5.8),
  • =$\mathcal{R}=\left\{{r_{a_{e},b_{e}}:e\in E_{1}}\right\}$ for $r_{a_{e},b_{e}}$ in Assumption 5.7(i), and
  • $\mathcal{C}$ the same bivariate copula densities for edges $e\in E_{j}$ with $j\geqslant 2$ as in (5.8).
Proposition 5.8 is foreshadowed by Theorem 3.4 and Example 3.5 in Li and Wu (2013), who consider D-vine copulas and who require stronger convergence properties.
5.3 Truncated X-vines
If the bivariate copula density $c_{a_{e},b_{e};D_{e}}$ associated with an edge $e\in E_{2}\cup\cdots\cup E_{d-1}$ is equal to the independence one, $c_{a_{e},b_{e};D_{e}}\equiv 1$, the corresponding factor drops out in (5.1) and the recursive formulas in (5.4) and (5.7) simplify to $R_{a_{e}|D_{e}\cup b_{e}}\left({x_{a_{e}}|\bm{x}_{D_{e}\cup b_{e}}}\right)=R_{a_{e}|D_{e}}\left({x_{a_{e}}|\bm{x}_{D_{e}}}\right)$ and $R_{b_{e}|D_{e}\cup a_{e}}\left({x_{b_{e}}|\bm{x}_{D_{e}\cup a_{e}}}\right)=R_{b_{e}|D_{e}}\left({x_{b_{e}}|\bm{x}_{D_{e}}}\right)$. Sparse X-vine specifications arise when $c_{a_{e},b_{e};D_{e}}\equiv 1$ for many edges $e\in E_{2}\cup\cdots\cup E_{d-1}$. Model selection of sparse vine copulas in high dimensions has been investigated in Müller and Czado (2019) and Nagler et al. (2019). The case where all pair copulas are equal to the independence one for all edges starting from a given tree is of practical importance (Brechmann et al., 2012; Brechmann and Joe, 2015).